A boy on a bridge throws a stone horizontally with a speed of 25 𝑚/𝑠, releasing the stone from a point 19.6 𝑚 above the surface of the river.  How far from a point directly below the boy will the stone strike the water?


1
Expert's answer
2021-09-23T08:32:49-0400

Let's first find the time taken by the stone to reach the river:


"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}}=\sqrt{\dfrac{2\cdot19.6}{9.8\ \dfrac{m}{s^2}}}=2.0\ s."

Finally, we can find the distance from the point directly below the boy to the point where the stone strikes the water:


"x=v_{0x}t=25\ \dfrac{m}{s}\cdot2.0\ s=50\ m."

Answer:

"x=50\ m."